Let X1, X2, X3 be uniform random variables on the interval (0, 1) with Cov(Xi , Xj ) = 1/24 for i, j = 1, 2, 3, i i != j. Calculate the variance of 2X1 + X2 − X3.

1
Expert's answer
2021-12-19T14:22:21-0500

Since "X_1, X_2, X_3"  are uniform random variables on the interval (0, 1), their probability distribution functions are,

 "f(X_1)=f(X_2)=f(X_3)= 1,\space 0\lt X_i\lt1, \space i=1,2,3"

"0, \space elsewhere"

"E(X_1)=\displaystyle\int^1_0 X_1dx_1={X_1^2\over 2}|^1_0={1\over2}"

"E(X_1^2)=\displaystyle\int^1_0 X_1^2dx_1={X_1^3\over3}|^1_0={1\over3}"

"Var(X_1)=E(X_1^2)-(E(X_1))^2={1\over3}-({1\over2})^2={1\over12}"


Since "X_2" and "X_3" have the same probability distribution function as "X_1", they both will have the same variance as "X_1", that is,

"Var(X_1)=Var(X_2)=Var(X_3)={1\over12}" and

"Cov(X_i,X_j)={1\over24}\space for\space i=1,2,3,\space j=1,2,3, \space i\not=j"

Now,

"Var(2X_1 + X_2 − X_3)=E[(2X_1 + X_2 − X_3)-(2\mu_1+\mu_2-\mu_3)]^2" , from definition

"=E[2(X_1-\mu_1)+(X_2+\mu_2)-(X_3-\mu_3)]^2"

"=4Var(X_1)+Var(X_2)+Var(X_3)+4Cov(X_1,X_2)-4Cov(X_1,X_3)-2Cov(X_2,X_3)"

"=4({1\over12})+{1\over12}+{1\over12}+4({1\over24})-4({1\over24})-2({1\over24})={6\over12}-{2\over24}={5\over12}"

Therefore,  the variance of 2X+ X2 − X3 is "{5\over12}".